# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ V Experiments
In this section, we present a comprehensive set of experiments designed to assess the performances of our SeqCRC approach to Object Detection (OD). These experiments include large-scale evaluations that result in a thorough benchmark, encompassing all matchings, losses, and prediction sets introduced in Section IV-A through IV-D.
@@
TABLE I: Comparison of set sizes and risks for all losses and prediction sets from Section IV, using DETR-50 as a reference model. The results are split into four categories: Confidence (Cnf.) and Localization (Loc.) losses, as well as Localization (Loc.) and Classification (Cls.) Prediction Sets (Pred. Sets). The risk levels are $\alpha^{\mathrm{cnf}}=0.02$ for the confidence task, $\alpha^{\mathrm{loc}}=0.05$ for localization, and $\alpha^{\mathrm{cls}}=0.05$ for classification. Thus, we apply the SeqCRC method at global level $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{loc}}+\alpha^{\mathrm{cls}}=0.1$ (see Corollary 1). For each setting, the results are shown using optimal choices for all remaining parameters.
@@
Loss Cnf. box_count_threshold 26.379 0.019 0.084
box_count_recall 18.868 0.018 0.085
Loc. thresholded 1.626 0.045 0.097
boxwise 1.578 0.049 0.098
pixelwise 1.056 0.051 0.100
Pred. Sets Loc. additive 1.094 0.051 0.099
multiplicative 1.056 0.051 0.100
Cls. aps 1.072 0.051 0.083
lac 1.044 0.051 0.083
@@ V-A Experimental Setup
In these experiments, we use two reference models: YOLOv8 [40] (in particular YOLOv8x [^14] https://github.com/ultralytics/ultralytics), a recent improvement in a long line of single-stage detectors including [38, 39], and DETR [43] (precisely, DETR-50151515https://github.com/facebookresearch/detr), a well-known transformer-based predictor. It is important to note that while all results here are based on these models, the method we present is model-agnostic and can be applied straightforwardly to virtually any other model (following our setting in Section II). The choice of these models emphasizes the broad applicability of our new framework, as well as the potential advantages or shortcomings of some families of models.
We employ the MS-COCO dataset [69], a standard benchmark in the field. For our calibration and inference procedures, we use the validation split of this dataset. This approach is common when an annotated dataset that is independent of training and model selection is unavailable. While potentially breaking Assumption 1, the practical availability constraints often make the validation split the best choice for conformal calibration and evaluation. The validation data is then split into two equal segments of $n=2500$ instances each, dedicated to calibration and inference respectively.
@@ V-A1 Methods
We conduct experiments on MS-COCO for both models (YOLO8 and DETR) and for all combinations of losses, distance functions (matchings), and prediction sets introduced in Section IV. For the matching “mix” referring to distance (10), we set $\tau=0.25$, providing a fair balance between localization and classification distances, according to preliminary experiments. For all experiments, the IoU threshold for the NMS is set to $0.5$. We focus on the setting of Corollary 1 that offers a joint guarantee on the localization and classification tasks (hence $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{cls}}+\alpha^{\mathrm{loc}}$ here). We consider two global error levels : $\alpha^{\mathrm{tot}}=0.1$ and $\alpha^{\mathrm{tot}}=0.2$. For the former we set $\alpha^{\mathrm{cnf}}=0.02$, $\alpha^{\mathrm{loc}}=0.05$ and $\alpha^{\mathrm{cls}}=0.05$ for confidence, localization and classification respectively. Similarly for the latter we set $\alpha^{\mathrm{cnf}}=0.03$, $\alpha^{\mathrm{loc}}=0.1$, $\alpha^{\mathrm{cls}}=0.1$. Finally, to mitigate the pathological second-step loss values towards $\lambda^{\mathrm{cnf}}=1$ (i.e. a null confidence threshold $1-\lambda^{\mathrm{cnf}}$) observed in Fig. 4, we filter out predictions with confidence lower than $10^{-3}$. This threshold is chosen independently of the data, for all experiments. (This thus corresponds to refining Steps 1 and 2 in Fig. 1, and does not break the guarantees of Theorem 2.) This leads to much lighter correction when applying the monotonization trick and therefore to smaller set sizes.
@@ V-A2 Metrics
For evaluation, we follow a common practice in Conformal Prediction, that of reporting the two metrics of Empirical Risk and Set Size. We examine their relationships with OD metrics in Appendix C-B. Throughout this section, we focus on the test data sequence $(X_{\mathrm{test},i},Y_{\mathrm{test},i})_{i=1}^{n_{\mathrm{test}}}$.
Firstly, the empirical risk (or simply risk thereafter) is the empirical estimation of the validity of the procedure, that is, the average loss measured empirically on test data. (In the case of a binary loss, it corresponds to $1$ minus the coverage, one of the main metrics in Conformal Prediction.) For each task $j\in\{\mathrm{cnf,loc,cls}\}$, the empirical risk that we measure is
@@
[…]
@@
[…]
@@
[…]
@@
GIoU 0.1 26.379 29.148 46.757
0.2 14.406 24.455 34.806
Hausdorff 0.1 26.379 1.056 44.271
0.2 14.406 1.000 23.466
LAC 0.1 26.379 15.239 1.044
0.2 23.974 8.314 0.669
Mix 0.1 26.379 1.396 8.709
0.2 23.974 1.043 0.906
TABLE III: Comparison of set sizes for all losses and prediction sets from Section IV for two models: DETR and YOLOX. The comparison is done for a global error level $\alpha^{\mathrm{tot}}=0.2$. The set size reported are the ones of the associated task in its part of the table.
@@
Loss Cnf. box_count_threshold 0.027 23.974 0.022 15.412
box_count_recall 0.026 14.406 0.031 10.138
Loc. thresholded 0.096 1.444 0.090 4.779
boxwise 0.105 1.322 0.096 3.935
pixelwise 0.098 1.000 0.100 2.562
Pred. Sets Loc. additive 0.098 1.000 0.101 2.957
multiplicative 0.093 1.006 0.100 2.562
Cls. aps 0.072 0.994 0.061 0.993
lac 0.101 0.669 0.106 0.545
V-B CODbase
To emphasize the reproducibility and reusability of our work, we introduce a comprehensive codebase, which we refer to as the “Conformal Object Detection Codebase,”161616 https://github.com/leoandeol/cods. or CODbase for short. This open-source repository includes several widely used object detectors, including, but not limited to, YOLO and DETR.
@@ V-B CODbase
The CODbase is implemented in Python and PyTorch, compatible with both CPU and GPU. Finally, the CODbase supports the integration of new procedures, prediction sets, and loss functions, enabling extensive benchmarks and further comparisons in future works.
@@ V-C Benchmark & Results
The benchmark consists of a large-scale study of all combinations of the losses, prediction sets, matching functions and more (see Section V-A). This results in the largest benchmark in Conformal Object Detection, with hundreds of individual runs. Key results are presented in several tables and figures in this section, while complete results are provided in Appendix C-A.
Impact of Component Choice We begin by examining Table I. For each loss and prediction set introduced in Section IV, we present the best experimental result (i.e., the reported experiment being the best combination of the remaining parameters). It does not include a section for both classification losses, and confidence prediction sets, as only one of each is introduced in Section IV. The table is organized as follows: it is split into two, with the upper part comparing different losses, and the lower part about prediction sets. These are then subdivided into tasks, where we compare losses or prediction sets specific to each task. For instance, the first line reads: for the box_count_threshold confidence loss, and the best remaining hyperparameters (other losses, matching, etc), the confidence set size is 26.379, with an empirical test risk of 0.019. The global (joint localization and classification) empirical risk is 0.084. The experiments are conducted with DETR-50, in the $\alpha^{\mathrm{tot}}=0.1$ setting. For virtually all tasks and global risk levels, the risk is controlled, i.e., below its corresponding risk level $\alpha^{\mathrm{tot}}$ or $\alpha^{j}$. A few experiments showcase a risk slightly above their respective $\alpha$, likely due to randomness from sampling. We also observe that the set sizes are sufficiently small and therefore practically informative for each loss. For example, in classification, when the classes in the set are semantically close, or induce a similar response, a set size of $3$ can be considered informative. For example, a set containing pedestrian, animal and cyclist for autonomous driving. The choice of loss function significantly affects the resulting set sizes, particularly for the losses presented here. The “pixelwise” loss, a relaxation of the previous two, produces smaller corrections at the cost of only slightly weaker guarantees (e.g., a guarantee on containing most of the ground truth rather than all of it), confirming results from [15]. Similarly, the relaxed confidence loss “box_count_recall” leads to the same conclusion.
@@
Matching Functions & Error Levels We now look at Table II. It compares, for each 4 proposed matching functions in Section IV-A, and two risk levels $\alpha^{\mathrm{tot}}$, the size metrics of the three tasks, for the best performing experiment in this setting (i.e. best losses and prediction sets for each $\alpha^{\mathrm{tot}}$ and matching function). We observe that the confident set size is constant for all matching functions for $\alpha^{\mathrm{tot}}=0.1$. This is due, in our setting, to the maximum of the three risks in Eq. (4) being always the confidence risk. Indeed, when $\alpha^{\mathrm{cnf}}$ is not too high, at least a box is selected on most images and therefore by definition, the localization and classification risks in this step are null (due to $\bar{\lambda}$). However, for $\alpha^{\mathrm{tot}}=0.2$, we observe two confident set sizes, which is due to the best experiment (in the sense of localization and classification set sizes) relying on different confidence losses depending on the matching. We further observe that localization and classification set sizes are respectively minimized with the “Hausdorff” and “LAC metrics”, while “Mix” produces an interesting trade-off of the two. However, the GIoU matching, while more OD-inspired, suffers from poor performance across the board, due to the lack of connection to the subsequent losses, that use the matching. Finally, in the “Mix” setting, a large jump in classification set size can be noted. Indeed, as observed in other experiments in Appendix C-A, the impact of a risk level $\alpha$ to its associated task’s resulting set size is not smooth, it can greatly vary for a small change in target risk.
Model-Agnosticism
We now look at Table III which compares set sizes and empirical risks for DETR and YOLOv8. It is built similarly to Table I, introduced previously. We observe that for both model, the risk is controlled for all experiments, that is below the corresponding risk level $\alpha^{\mathrm{tot}}$ or $\alpha^{j}$. [^17] Using APS results in a looser bound, as discussed in [65] This result is a direct consequence of the model-agnosticism of conformal prediction and therefore of SeqCRC – it is applicable regardless of the performance or size of the model. However, the set size depends on the quality of the underlying model. Furthermore, there is a trade-off in terms of number, and size of (conformal) bounding boxes. We observe here that using YOLOv8 tends to lead to much fewer selected bounding boxes, but with much larger corrections in localization: with the same hyperparameters used for calibration, the resulting trade-off differs.
@@ V-D Discussion
Risk Control We have observed that in all experiments, the risk is controlled (or, at worst, very close to the target), empirically verifying the guarantees presented in Theorem 2. However, it appears that the global risk guaranteed by Corollary 1 can be unnecessarily conservative, as observed in Table I and in Appendix C-A. The correction applied to obtain this global guarantee can therefore be a future point of refinement.
Guarantee strength – set size Tradeoff A shown in Table I and Table III, the further relaxed losses, as the “pixelwise” in localization, offer much smaller prediction sets. This however comes at a cost of a slightly weaker guarantee. In that case again, the two other losses provide a guarantee on the whole object being covered, while the “pixelwise”, only on the proportion of the area that is covered (i.e., missing a few pixels of a ground truth is not too punitive). We argue this approach is generally sensible and beneficial, but in some critical applications, a guarantee that the entire ground truth is contained is necessary, such as in [70].
Strengths and Limitations in Practice Turning to a more qualitative analysis, we make further observations. There are many examples of successful conformal predictions such as on Fig. 7, where the right amount of object is output, entirely contained by the slightly-corrected conformal boxes, and with a correct, singleton, classification prediction set. However, there are also shortcomings of our approach, and we focus on an example in Fig. 8. There, while the two zebras have rather satisfactory conformal predictions, there is an additional large prediction including both animals, and not linked to any ground truth. This allows us to raise two important points.
First, our approach only provides Recall guarantees, and therefore does not treat false positives as an issue. In fact, due to precision being a strongly non-monotone function of our parameters, monotonizing it while using our approach generally fails (due to the monotonized risk being too large at its lowest value).
Secondly, we have observed that the annotation in MS-COCO is very variable, and one such example, is when there are multiple occurrences of an object. Often, some individual objects are annotated, but also, there is an additional box, including all these objects, and many non annotated ones. Several such examples are presented in Appendix C-B. The model being trained on such data, repeats this pattern on some test images when not expected, and sometimes not when it is.
Finally, we recall that conformal prediction brings a guarantee on the loss of the prediction with regards to the ground truth. If the ground truth is incorrect, or unpredictable (due to inconsistencies in the annotation choices or process), and if the model does not make those errors, it will lead to arbitrarily large corrections necessary to enforce low risk on the conformal predictions, resulting in practically unusable sets.
Figure 8: Example of an artifact in conformal predictions. We observe that beyond the two correctly detected zebras, is an additional, larger, third prediction that includes both animals, but does not correspond to a ground truth. We argue this is due to artifacts in the annotation of MS-COCO, and therefore due to the pretrained model.

### v2

@@ V Experiments
In this section, we present a comprehensive set of experiments designed to assess the performances of our SeqCRC approach to Object Detection (OD). These experiments include large-scale evaluations that result in a thorough benchmark, encompassing all matchings, losses, and prediction sets introduced in Sections IV-A through IV-D.
@@
TABLE I: Comparison of set sizes and risks for all losses and prediction sets from Section IV, using DETR-101 as a reference model. The results are split into four categories: Confidence (Cnf.) and Localization (Loc.) losses, as well as Localization (Loc.) and Classification (Cls.) Prediction Sets (Pred. Sets). The risk levels are $\alpha^{\mathrm{cnf}}=0.02$ for the confidence task, $\alpha^{\mathrm{loc}}=0.05$ for localization, and $\alpha^{\mathrm{cls}}=0.05$ for classification. Thus, we apply the SeqCRC method at global level $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{loc}}+\alpha^{\mathrm{cls}}=0.1$ (see Corollary 1). For each setting, the results are shown using optimal choices for all remaining parameters.
@@
Loss
Cnf.
box_count_threshold 25.588 0.022 0.086
box_count_recall 17.778 0.019 0.085
Loc.
thresholded 1.552 0.046 0.097
boxwise 1.504 0.049 0.097
pixelwise 1.043 0.047 0.096
Pred. Sets
Loc.
additive 1.047 0.052 0.100
multiplicative 1.043 0.047 0.096
Cls.
aps 1.007 0.050 0.082
lac 0.994 0.051 0.087
@@ V-A Experimental Setup
In these experiments, we use two reference models: YOLOv8 [40] (more precisely, YOLOv8xhttps://github.com/ultralytics/ultralytics, 68M parameters), a recent improvement in a long line of single-stage detectors including [38, 39], and DETR [43] (DETR-101https://github.com/facebookresearch/detr, 60M parameters), a well-known transformer-based predictor. It is important to note that while all results here are based on these models, the method we present is model-agnostic and can be applied straightforwardly to virtually any other model (following our setting in Section II). The choice of these models emphasizes the broad applicability of our new framework, as well as the potential advantages or shortcomings of some families of models.
We employ the MS-COCO dataset [69], a standard benchmark in the field. For our calibration and inference procedures, we use the validation split of this dataset. This approach is common when an annotated dataset that is independent of training and model selection is unavailable [53, 22]. While potentially breaking Assumption 1, the practical availability constraints often make the validation split the best choice for conformal calibration and evaluation. The validation data is then split into two equal segments of $n=2500$ instances each, dedicated to calibration and inference respectively.
@@ V-A1 Methods
We conduct experiments on MS-COCO for both models (YOLO8 and DETR) and for all combinations of losses, distance functions (matchings), and prediction sets introduced in Section IV. For the matching “mix” referring to distance (10), we set $\tau=0.25$, providing a fair balance between localization and classification distances, according to preliminary experiments. For all experiments, the IoU threshold for the NMS is set to $0.5$. We focus on the setting of Corollary 1 that offers a joint guarantee on the localization and classification tasks (hence $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{cls}}+\alpha^{\mathrm{loc}}$ here). We consider two global error levels : $\alpha^{\mathrm{tot}}=0.1$ and $\alpha^{\mathrm{tot}}=0.2$. For the former we set $\alpha^{\mathrm{cnf}}=0.02$, $\alpha^{\mathrm{loc}}=0.05$ and $\alpha^{\mathrm{cls}}=0.05$ for confidence, localization and classification respectively. Similarly for the latter we set $\alpha^{\mathrm{cnf}}=0.03$, $\alpha^{\mathrm{loc}}=0.1$, $\alpha^{\mathrm{cls}}=0.1$. Finally, to mitigate the pathological second-step loss values towards $\lambda^{\mathrm{cnf}}=1$ (i.e. a null confidence threshold $1-\lambda^{\mathrm{cnf}}$) as illustrated in Fig. 4, we filter out predictions with confidence lower than $10^{-3}$. This threshold is chosen independently of the data, for all experiments. (This thus corresponds to refining Steps 1 and 2 in Fig. 1, and does not break the guarantees of Theorem 2.) This leads to much lighter correction when applying the monotonization trick and therefore to smaller set sizes.
@@ V-A2 Metrics
For evaluation, we follow a common practice in Conformal Prediction, that of reporting the two metrics of Empirical Risk and Set Size. We examine their relationships with OD metrics in Appendix E-B. Throughout this section, we focus on the test data sequence $(X_{\mathrm{test},i},Y_{\mathrm{test},i})_{i=1}^{n_{\mathrm{test}}}$.
The empirical risk (or simply risk) measures the average loss on test data. For binary losses, this equals $1$ minus the coverage, a standard metric in Conformal Prediction. (In the case of a binary loss, it corresponds to $1$ minus the coverage, one of the main metrics in Conformal Prediction.) For each task $j\in\{\mathrm{cnf,loc,cls}\}$, the empirical risk that we measure is
@@
$\text{Confidence Set Size:}\quad\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}_{+}}(X_{\mathrm{test},i})\rvert$.
@@
$\text{Localization Set Size:}\quad\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\frac{1}{n_{\mathrm{test},i}}\sum_{\hat{b}_{k},\hat{b}_{k}^{\lambda^{\mathrm{loc}}_{+}}}\sqrt{\frac{\mathrm{area}(\hat{b}_{k}^{\lambda^{\mathrm{loc}}_{+}})}{\mathrm{area}(\hat{b}_{k})}},$
@@
$\text{Classification Set Size:}\quad\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\frac{1}{n_{\mathrm{test},i}}\sum_{\hat{c}_{k}^{\lambda^{\mathrm{cls}}_{+}}}\lvert\hat{c}_{k}^{\lambda^{\mathrm{cls}}_{+}}\rvert$,
@@
GIoU 0.1 17.778 28.241 44.471
0.2 14.046 23.690 32.335
Hausdorff 0.1 25.588 1.043 41.846
0.2 14.046 0.999 22.035
LAC 0.1 25.588 14.147 0.994
0.2 22.657 7.786 0.653
Mix 0.1 25.588 1.334 8.228
0.2 22.657 1.018 0.931
TABLE III: Comparison of set sizes for all losses and prediction sets from Section IV for two models: DETR-101 and YOLOv8x. The comparison is done for a global error level $\alpha^{\mathrm{tot}}=0.1$. The set size reported are the ones of the associated task in its part of the table.
@@
Loss
Cnf.
box_count_threshold 0.022 25.588 0.012 18.855
box_count_recall 0.019 17.778 0.019 11.710
Loc.
thresholded 0.046 1.552 0.047 5.451
boxwise 0.049 1.504 0.046 5.225
pixelwise 0.047 1.043 0.049 3.867
Pred. Sets
Loc.
additive 0.052 1.047 0.049 3.867
multiplicative 0.047 1.043 0.052 4.181
Cls.
aps 0.050 1.007 0.043 0.996
lac 0.051 0.994 0.049 0.717
V-B COD Toolkit
To emphasize the reproducibility and reusability of our work, we introduce a comprehensive toolkit, which we refer to as the “Conformal Object Detection Toolkit,”https://github.com/leoandeol/cods. or COD for short. This open-source repository includes several widely used object detectors, including, but not limited to, YOLO and DETR.
@@ V-B COD Toolkit
The COD toolkit is implemented in Python and PyTorch, compatible with both CPU and GPU. Finally, it supports the integration of new procedures, prediction sets, and loss functions, enabling extensive benchmarks and further comparisons in future works.
@@ V-C Benchmark & Results
The benchmark consists of a large-scale study of all combinations of the losses, prediction sets, matching functions and more (see Section V-A). This results in the largest benchmark in Conformal Object Detection, with hundreds of individual runs. Key results are presented in several tables and figures in this section, while complete results are provided in Appendix E-A.
Impact of Component Choice. We begin by examining Table I. For each loss and prediction set introduced in Section IV, we present the best experimental result (i.e., the reported experiment being the best combination of the remaining parameters). It does not include a section for both classification losses, and confidence prediction sets, as only one of each is introduced in Section IV. The table is organized as follows: it is split into two, with the upper part comparing different losses, and the lower part about prediction sets. These are then subdivided into tasks, where we compare losses or prediction sets specific to each task. For instance, the first line reads: for the box_count_threshold confidence loss, and the best remaining hyperparameters (other losses, matching, etc), the confidence set size is 25.588, with an empirical test risk of 0.022. The global (joint localization and classification) empirical risk is 0.086. The experiments are conducted with DETR-101, in the $\alpha^{\mathrm{tot}}=0.1$ setting. For virtually all tasks and global risk levels, the risk is controlled, i.e., below its corresponding risk level $\alpha^{\mathrm{tot}}$ or $\alpha^{j}$. A few experiments showcase a risk slightly above their respective $\alpha$, likely due to randomness from sampling. We also observe that the set sizes are sufficiently small and therefore practically informative for each loss. For example, in classification, when the classes in the set are semantically close, or induce a similar response, a set size of $3$ can be considered informative. For example, a set containing pedestrian, animal and cyclist for autonomous driving. The choice of loss function significantly affects the resulting set sizes, particularly for the losses presented here. The “pixelwise” loss, a relaxation of the previous two, produces smaller corrections at the cost of only slightly weaker guarantees (e.g., a guarantee on containing most of the ground truth rather than all of it), confirming results from [15]. Similarly, the relaxed confidence loss “box_count_recall” leads to the same conclusion.
@@
Matching Functions & Error Levels. We now look at Table II. It compares, for each 4 proposed matching functions in Section IV-A, and two risk levels $\alpha^{\mathrm{tot}}$, the size metrics of the three tasks, for the best performing experiment in this setting (i.e. best losses and prediction sets for each $\alpha^{\mathrm{tot}}$ and matching function). We observe that the confident set size is constant for all matching functions for $\alpha^{\mathrm{tot}}=0.1$. This is due, in our setting, to the maximum of the three risks in Eq. (4) being always the confidence risk. Indeed, when $\alpha^{\mathrm{cnf}}$ is not too high, at least a box is selected on most images and therefore by definition, the localization and classification risks in this step are null (due to $\bar{\lambda}$). However, for $\alpha^{\mathrm{tot}}=0.2$, we observe two confident set sizes, which is due to the best experiment (in the sense of localization and classification set sizes) relying on different confidence losses depending on the matching. We further observe that localization and classification set sizes are respectively minimized with the “Hausdorff” and “LAC metrics”, while “Mix” produces an interesting trade-off of the two. However, the GIoU matching, while more OD-inspired, suffers from poor performance across the board, due to the lack of connection to the subsequent losses, that use the matching. Finally, in the “Mix” setting, a large jump in classification set size can be noted. Indeed, as observed in other experiments in Appendix E-A, the impact of a risk level $\alpha$ to its associated task’s resulting set size is not smooth, it can greatly vary for a small change in target risk.
Model-Agnosticism. We now look at Table III which compares set sizes and empirical risks for DETR-101 and YOLOv8x. It is built similarly to Table I, introduced previously. We observe that for both model, the risk is controlled for all experiments, that is below the corresponding risk level $\alpha^{\mathrm{tot}}$ or $\alpha^{j}$.Using APS results in a looser bound, as discussed in [65]. This result is a direct consequence of the model-agnosticism of conformal prediction and therefore of SeqCRC – it is applicable regardless of the performance or size of the model. However, the set size depends on the quality of the underlying model. Furthermore, there is a trade-off in terms of number, and size of (conformal) bounding boxes. We observe here that using YOLOv8 tends to lead to much fewer selected bounding boxes, but with much larger corrections in localization: with the same hyperparameters used for calibration, the resulting trade-off differs.
@@ V-D Discussion
Risk Control. In all our experiments, the risk meets or is very close to the target value, which empirically confirms Theorem 2. However, the global risk from Corollary 1 is often unnecessarily conservative (see Table I and Appendix E-A), suggesting that its correction could be refined.
Tradeoff: Guarantee Strength versus Set Size. As shown in Tables I and III, relaxed losses such as the “pixelwise” localization loss yield much smaller prediction sets, but at the cost of weaker guarantees. Unlike the stricter losses, which ensure full object coverage, the pixelwise variant only guarantees that a proportion of the ground-truth area is covered (i.e., missing a few pixels is not heavily penalized). This tradeoff is often acceptable, but in some safety-critical applications (e.g., [70]) full-object guarantees remain necessary.
Strengths and Limitations in Practice. Qualitatively, many examples (e.g., Fig. 6) show accurate object counts, predictions that are fully contained in conformal boxes, and correct singleton classification sets. Limitations, however, also appear: in Fig. 8, while the two zebras are well predicted, an additional large box includes both animals without ground-truth support, which reveals two issues.
First, our approach guarantees a large enough recall but only empirically limits false positives (via the confidence thresholding step). In fact, precision is a non-monotone function of its parameters, and monotonizing it using our approach could lead to too large risks for a solution to exist.
Second, MS-COCO annotations are inconsistent: individual objects may be labeled alongside a larger enclosing box, and sometimes include unlabeled instances (see Appendix E-B). Models trained on such data reproduce this variability, generating group boxes unpredictably.
Finally, we recall that conformal prediction brings a guarantee on the loss of the prediction with regards to the ground truth. If the ground truth is incorrect or inconsistent (e.g., due to annotation choices or labeling errors), and the model in fact predicts the object more accurately than the labels, then conformal prediction must impose very large corrections in order to achieve small risk. These corrections could inflate the prediction sets to the point of becoming practically unusable.
Figure 8: Example of an annotation/model artifact. Beyond the two correctly detected zebras, an additional (larger) prediction includes both animals, but does not correspond to any ground truth. We argue this is due to artifacts in the annotation of MS-COCO, which are reproduced by the pretrained model.